OpenStudy (anonymous):

(8-3√6)(7+2√6) Answers: 20-37√6 56-6√6 56-11√6 20-5√6

OpenStudy (saifoo.khan):

D for Duck

OpenStudy (saifoo.khan):

\[\Huge \begin{array}l\color{red}{\text{D}}\color{orange}{\text{D}}\color{#9c9a2e}{\text{D}}\color{green}{\text{D}}\color{blue}{\text{D}}\color{purple}{\text{D}}\color{purple}{\text{D}}\color{red}{\text{D}}\color{orange}{\text{D}}\color{#9c9a2e}{\text{D}}\color{green}{\text{D}}\color{blue}{\text{D}}\color{purple}{\text{D}}\color{purple}{\text{D}}\color{red}{\text{D}}\color{orange}{\text{D}}\color{#9c9a2e}{\text{D}}\color{green}{\text{D}}\color{blue}{\text{D}}\color{purple}{\text{D}}\color{purple}{\text{D}}\color{red}{\text{D}}\color{orange}{\text{D}}\color{#9c9a2e}{\text{D}}\color{green}{\text{D}}\color{blue}{\text{D}}\color{purple}{\text{D}}\color{purple}{\text{D}}\color{red}{\text{D}}\color{orange}{\text{D}}\color{#9c9a2e}{\text{D}}\color{green}{\text{D}}\color{blue}{\text{D}}\color{purple}{\text{D}}\color{purple}{\text{d}}\color{red}{\text{}}\end{array}\]

OpenStudy (anonymous):

D FOR APPLE

OpenStudy (anonymous):

20-5√6

OpenStudy (saifoo.khan):

\[\Huge \color\purple {20-5\sqrt{6}}\]

OpenStudy (angela210793):

56+16sqr6-21sqr6-36=20-5sqr6